1)find the perimeter of a circle with radius = 21cm


2)find the area and perimeter of akite with lengths 20cm and 14cm


1
Expert's answer
2022-01-13T16:10:05-0500

1)


"Perimeter=2\pi r=2\pi (21cm)=42\pi cm\approx"

"\approx131.947cm"

2)


"Perimeter=2(20cm+14cm)=68 cm"

I cannot find the area of a kite because


"Area=\dfrac{d_1d_2}{2},"

where "d_1, d_2" are the diagonals of the kite, but the diagonals are unknown.